OpenStudy (anonymous):

3x8 – 7x8 –4x16 –21x8 –4x8 10x8

OpenStudy (kinggeorge):

-4x8\[3\cdot8-7\cdot8=(3-7)\cdot8=-4\cdot8\]

OpenStudy (anonymous):

-4x8 3⋅8−7⋅8=(3−7)⋅8=−4⋅8 I think...